OpenStudy (anonymous):

how do I solve 3x+-4=x+4

OpenStudy (amoodarya):

do you mean ?\[3x\pm4=x+4 \]

OpenStudy (anonymous):

yes

Atsie (atsie):

Firstly you need to reorder everything -4 + 3x = x + 4 and -4 + 3x = 4 + x

OpenStudy (amoodarya):

do it for two case : 1 3x+4=x+4 2 3x-4=x+4

Atsie (atsie):

then solve for the variable x by move all your terms containing x to the left, and all other terms to the right.

Atsie (atsie):

so it should look like this... \[-4 + 3x + -1x = 4 + x + -1x\]

OpenStudy (anonymous):

is a step to solving for x

OpenStudy (amoodarya):

−4+3x+(−1x)=4+x+(−1x) ? if you mean it ,you'r right

Atsie (atsie):

@djHunter You following all this?

OpenStudy (amoodarya):

let me see your work